Page tree

Anuket Project

Skip to end of metadata
Go to start of metadata

The weekly meeting takes place on Tuesdays at 17:00 Irish Standard time (UTC)

Meeting is held on zoom: https://zoom.us/j/92950950555?pwd=K2tWZWovWXJWTmNuZHIxNGlZWkpHQT09 and IRC #opnfv-meeting.


Standing agenda items

  • TSC updates
  • CI status
  • Release updates
  • Action item review


Next meeting

The next meeting takes place on , at 16h00 UTC (17h00 Ireland/Dublin).

Agenda

01-02-2022

  • TSC updates
  • CI status
  • Release updates
  • Action item review

2022 Meetings

25-Jan-2022

18-Jan-2022

  • Release updates
    • Moselle is kicking off
  • CI status
    • All are green!
  • TSC updates
  • Standards updates

11-Jan-2022

  • cancelled

04-Jan-2022

  • Release updates
    • Tasks complete for Lakelse M5
  • CI status
    • All green, and Lakelse jobs are up
  • TSC updates
  • Standards updates

2021 Meetings

14-December-2021

  • Lakelse updates
    • exception approved for barometer m5 until December 17th
  • CI status
    • All green!
  • TSC updates
  • Standards updates

07-December-2021

30-November-2021

  • Lakelse Updates
    • TSC approved M4
  • CI status
    • Everything is passing!
  • TSC updates
  • Standards updates
    • The TST008 update is that the ETSI Editors contacted Al today to say that their editing process (on the approved text) has begun.

23-November-2021

  • Lakelse Updates
    • Release will be early next year
  • CI status
    • Everything is passing!
  • TSC updates
  • Standards updates
    • No comment; remote consensus process closed; Document number updated to NFV(21)000243, and it's now considered "done"; expected actions from editors to update style and references;

16-November-2021

  • Lakelse Updates
    • M4 is up today, which applies to specification projects
    • There was a discussion on moving the release schedule so Lakelse finishes later
  • CI status
    • Everything is passing
  • TSC updates
  • Standards updates

09-November-2021

02-November-2021

26-October-2021

19-October-2021

12-October-2021

05-October-2021

  • Lakelse update
    • M3 is October 19th
  • CI status
    • Everything is passing
  • TSC Updates
    • Nominations are open for the 2022 TSC memberships
  • Standards updates
    • No updates; latest meeting was post-poned
  • Committer clean-up
    • Process for removing committers:- either a committer steps down voluntarily or can be removed by a 2/3 majority vote.
    • Action(Emma) Reach out to Calin about stepping down

28-September-2021

  • Lakelse update
  • CI status
  • TSC Updates
    • Nominations are open for the 2022 TSC memberships
  • Standards updates
    • A complete working group review of the document is happening, then it will be submitted for full approval.
  • Committer clean-up
    • Emma did not check the existing process
    • Emma e-mailed Aaron about stepping down as a committer - he agreed
    • Al e-mailed Maryam about stepping down as committer - she sadly agreed
    • Action(Emma): Check the existing process for removing committers;
    • Action(Emma) Reach out to Calin about stepping down

21-September-2021

31-August-2021

  • Lakelse updates
    • Today is M2 - scope freeze
  • CI status
  • Standards updates
    • The ETSI-NFV test working group accepted the contribution on storage!
  • AoB

24-August-2021

  • Lakelse updates
    • Need to add Lakelse labels to Jira items
  • CI status
  • Standards updates
    • Updated draft of TST-008 to include storage metrics
  • AoB

17-August-2021

  • Lakelse updates:
    • M2 (scope freeze) is  
    • Discussions/triaging/scoping of RA reqs into to Jira items is taking place in the Release meeting for the next 2 weeks.
  • Ci status
    • No status change
  • Standards updates
    • no status change
  • AoB

10-August-2021

  • Lakelse updates:
    • M2 (scope freeze) is  
    • Discussions/triaging/scoping of RA reqs into to Jira items is taking place in the Release meeting for the next 3 weeks.
  • Ci status
    • Everything looks fine (some older jobs are failing, but that's nothing to worry about)
  • Standards updates
    • No updates
  • AoB

03-August-2021

  • Lakelse updates:
    • M2 (scope freeze) is  
    • Discussions/triaging/scoping of RA reqs into to Jira items is taking place in the Release meeting for the next 4 weeks.
  • Ci status
    • barometer-daily-master is working again!
  • Thoth, possible demo/collab, discussion on the weekly tech discuss and weekly AI/ML meeting.
    • Thoth project got approved at this week's TSC!
  • AoB

13-Jul-2021

  • Announcement: Meetings are cancelled the next two weeks for vacation time
  • Lakelse updates:
    • M2 (scope freeze) is  
  • Ci status: barometer-daily-master is mostly working; there's an issue in uploading the artifacts, I've logged a ticket to get it investigated.
  • Gitlab transition:
    • ETA is mid-August
  • AI/ML project, possible demo/collab, discussion on the weekly tech discuss and weekly AI/ML meeting.
  • AoB

06-Jul-2021

Older meetings






  • No labels

1 Comment

  1. I mentioned this during the Dec 18 VSPERF meeting. PTL @Sridhar Rao has been collecting a list and he will communicate to me/Emma in time for next Barometer meeting (2020).